OpenStudy (anonymous):

verify. cos^2x-sin^2x/cos^2x+sinxcosx=1/tanx

OpenStudy (shubhamsrg):

take cos x common from denon and see if it works ?

OpenStudy (lgbasallote):

would be wonderful if this was rewritten to show how the fractions look like

OpenStudy (anonymous):

\[\cos ^{2}x-\sin ^{2}x \div \cos ^{2}x + sinxcosx = 1/tanx\]

OpenStudy (lgbasallote):

there's also a button in the equation editor for fractions...you might want to use that

OpenStudy (anonymous):

nevermind I figured it out :) thanks though!

OpenStudy (anonymous):

\[CosxTan ^{3}x = SinxTan ^{2}x\]